# REPLAY: Modeling Time-Varying Temporal Regularities of Human Mobility for Location Prediction over Sparse Trajectories

### v1

@@ 4.2 Location Prediction Performance
Table 1 shows the location prediction performance of different methods. First, we observe that REPLAY achieves the best performance in most cases, yielding 7.7% and 10.9% improvement over the best-performing baselines on Gowalla and Foursquare, respectively. Moreover, REPLAY significantly outperforms the best baselines in Acc@1 (by 23.4% and 25.8% on Gowalla and Foursquare, respectively). The only exception is on Acc@10 where Graph-Flashback achieves slightly better performance than REPLAY; note that beyond user mobility trajectories, Graph-Flashback further uses user social relationships as input.
@@
We also highlight in Table 1 the nine methods that make predictions using query time information in different formats (e.g., a query timestamp, a query time interval, or a query time threshold). However, we did not observe a consistent improvement of these methods over others that make predictions without using query time information. This is due to the fact that the prediction performance of these baseline methods is dominated by their distinct model structures, rather than the benefit of adding query time information.
Finally, we also conduct a systematic ablation study of REPLAY in Appendix D to verify the usefulness of both our smoothed timestamp embeddings (with 20.6% and 7.5% improvement on Gowalla and Foursquare, respectively) and the query time (with 18.0% and 11.1% improvement on Gowalla and Foursquare, respectively) in location prediction.
Table 3: Impact of different time granularities and scales on location prediction performance
Scale Granu- larity Number of Timestamps Gowalla Foursquare
Acc@1 Acc@5 Acc@10 MRR
Time per
epoch
Acc@1 Acc@5 Acc@10 MRR
Time per
epoch
Day Minute 1440 0.1278 0.2918 0.3610 0.2060 41.35s 0.2801 0.5723 0.6468 0.4121 122.14s
Hour 24 0.1777 0.3411 0.4063 0.2555 39.35s 0.3483 0.5927 0.6622 0.4601 101.85s
Weekday& Weekend Minute 2880 0.1804 0.3423 0.4090 0.2577 44.98s 0.3530 0.5938 0.6631 0.4633 142.74s
Hour 48 0.1821 0.3442 0.4097 0.2594 37.69s 0.3505 0.5941 0.6632 0.4618 102.33s
Week Minute 10080 0.1756 0.3276 0.3891 0.2485 72.41s 0.3532 0.5940 0.6638 0.4636 264.00s
Hour 168 0.1866 0.3476 0.4124 0.2635 44.20s 0.3529 0.5953 0.6648 0.4638 103.49s
4.3 Different Time Granularities and Scales
We study the impact of different time granularities (minute or hour) and time scales (day, weekday&weekend or week) in our smoothed timestamp embeddings. Note that for the time scale weekday&weekend, we define two sets of daily cyclical timestamps for weekday and weekend, respectively; the timestamp distance in each set is computed independently. Table 3 shows the different settings and their results.
First, we observe that the week scale consistently outperforms the day scale by 2.89% and 0.74% (on the hour granularity) on Gowalla and Foursquare, respectively. Because the former can capture the different temporal regularities across different days in a week, while the latter cannot. Moreover, compared to the weekday&weekend scale that captures the different regularities on weekdays and weekends, the week scale still achieves better performance (with 1.43% and 0.4% improvement on Gowalla and Foursquare, respectively, under the hour granularity). Because the week scale can further capture the regularity differences between different weekdays (also between different days on the weekend).
Second, we observe that the hour granularity is generally better than the minute granularity, as the minute granularity is often over-specific to model the temporal regularity of user mobility. One exception is on the week scale on Foursquare where we observe comparable results; however, in this case, the minute granularity shows significantly worse runtime performance (more than twice the training time as shown in Table 3) than the hour granularity, as it needs to learn much more embeddings for minute-in-week timestamps (10,080 in total) than for hour-in-week timestamps (168 in total). Moreover, we also conducted experiments on multi-granularity combining both hour-in-day and day-in-week timestamp embeddings in Appendix D; the results show 3.4% and 0.6% improvement of our REPLAY over this multi-granularity setting, on Gowalla and Foursquare, respectively. Therefore, we use hour-in-week timestamps in REPLAY.
(a) Hours in week
(b) Hours on weekdays
(c) Hours on weekends
Figure 4: The learnt bandwidths and the corresponding performance across different timestamps on Foursquare.
4.4 Revealing the Time-Varying Temporal Regularities
We reveal the time-vary temporal regularities of user mobility using the learnt bandwidths from REPLAY. Figure 5(a) shows the values of the learnt bandwidths for the 168 hour-in-week timestamps and the corresponding location prediction performance in MRR for each timestamp on Foursquare (similar results on Gowalla as shown in Appendix E).
[…]
Second, we also observe a clear daily pattern in location prediction performance, which, interestingly, shows an inverse trend compared to the bandwidths. This is due to the fact that the weaker regularity of user mobility at a certain timestamp (in the nighttime or on weekends, for example) makes it harder for location prediction; subsequently, the learnt bandwidth has a larger value, as the query timestamp is less confident for location prediction and thus resorts more to its neighboring timestamps for prediction.
Table 4: Location prediction performance and the corresponding average bandwidth $\bar{\sigma}$ in different time periods.
Prediction time period Gowalla Foursquare
MRR
$\bar{\sigma}$
MRR
$\bar{\sigma}$
Daytime(6:00-18:00)
0.2767 0.85 0.4856 1.76
Nighttime(18:00-6:00)
0.2362 0.88 0.4302 2.00
Weekday 0.2922 0.82 0.4841 1.84
Weekend 0.1898 0.98 0.3758 2.00
Finally, we quantitatively verify our motivating example where the strength of the temporal regularities varies across typical time periods (daytime/nighttime and weekday/weekend). Table 4 shows the average MRR and learnt bandwidths in these periods. We see the daytime prediction performance is 15.0% higher (7.5% lower bandwidth) than the nighttime, while the weekday prediction performance is 41.4% higher (11.9% lower bandwidth) than the weekend.
5 Conclusion
In this paper, we propose REPLAY, a general RNN architecture capturing the time-varying temporal regularities of user mobility for location prediction. It seamlessly incorporates smoothed timestamp embeddings with learnable bandwidths into a flashback mechanism. The timestamp-specific learnable bandwidths can automatically learn to adapt to the temporal regularities of different strengths across different timestamps. We conduct a thorough evaluation using two real-world LBSN datasets. Results show that REPLAY outperforms a sizable collection of state-of-the-art location prediction methods, with 7.7%-10.9% improvement over best-performing baselines, and the learnt bandwidths reveal interesting patterns of the time-varying temporal regularities.
In future work, we plan to study the temporal regularities varying across different user groups and POI categories.
Appendix A The details of our Model Training
The training process of REPLAY follows the traditional backpropagation-through-time training process of recurrent neural networks. For each prediction, we compute the cross-entropy loss between the probability distribution of POIs output by REPLAY and the ground truth POIs, which are minimized using Adam optimizer. The loss function is defined as follows:
where $U$ is the number of users; $N_{j}$ is the sequence length of $j$-th user; $y_{i+1,j}$ is the one-hot vector of ground truth POI for ($i$+$1$)-th check-in of $j$-th user. $p(\hat{y}_{i+1,j}|\mathcal{H}_{i},s_{k},u_{j})$ is the predicted probability distribution given the current aggregated hidden state $\mathcal{H}_{i}$, smoothed query timestamp $s_{k}$ (transformed from time $t_{i+1}$), and user embedding $u_{j}$.
Appendix B The details of baselines
The details of baselines and the hyperparameters as follow:
  • User Preference-based Methods. WRMF is a matrix factorization based recommendation technique designed explicitly for implicit feedback such as historical user behavior, where users do not explicitly express their preference; it fits well our check-in datasets that contain historical user presence at POIs. BPR is a ranking-based recommendation technique for implicit feedback, which learns user preferences by minimizing a pairwise ranking loss; it uses a bootstrap sampling technique to learn from both positive feedback and non-observed feedback to alleviate data sparsity issue. We use the default parameter settings of LibRec [^2] https://github.com/guoguibing/librec for these two methods.
  • Feature-based methods: Most Frequent Time (MFT) ranks a POI according to a user’s historical check-in count at a POI and at a specific time slot in the training dataset; we define the timeslot as 24 hours in a typical day as suggested by paper. LBSN2Vec is an automatic feature learning technique designed specifically for check-in data; it learns user, time and POI feature vectors from an LBSN hypergraph, and ranks a POI according to its similarities with user and time in the feature space for location prediction tasks; we set the number of negative samples, random walk window size, and the mobility data ratio to {10, 10, 1} in our experiments.
  • Markov-Chain-based Methods: FPMC estimates a personalized transition matrix of POIs in user mobility trajectories using matrix factorization techniques. We set the number of negative samples and the dimension of factorization to {10, 32}. PRME extends FPMC to location prediction problems by further considering spatial constraints, and learns user and POI embeddings to capture the personalized POI transition patterns. We set the threshold, component weight, and hidden state size to {360, 0.2, 20}. TribeFlow learns a mixture model to capture the semi-Markov transition probability matrix over latent environments for predicting user trajectories. We set the number of cores, number of transitions, number of iterations, and the number of batches to {20, 0.3, 2000, 20}, and use ECCDF-based kernel introduced in paper.
  • Basic RNNs: RNN is a vanilla Recurrent Neural Network architecture, capturing sequential patterns from user mobility trajectories for location prediction. LSTM (Long Short-Term Memory) is capable of learning from both short- and long-term dependency in sequences using a memory cell controlled by three multiplicative gates including an input gate, an output gate and a forget gate. GRU (Gated Recurrent Unit) captures long-term dependency by controlling information flow with an update gate and a reset gate. We set the hidden state size to 10 for all three RNN architectures and all the datasets.
  • Spatiotemporal Sequence Models: DeepMove adds an attention mechanism to GRU for location prediction over sparse mobility trajectories. We set the size of all embeddings and the hidden state size to {50, 10}, and use average history mode and LSTM as suggested by paper. STRNN uses customized transition matrices parameterized by the spatiotemporal distances between check-ins within a time window in RNNs. STGN adds additional gates controlled by the spatiotemporal distances between successive check-ins to LSTM. STGCN is an extension of STGN with coupled input and forget gates for improved efficiency. We set the hidden state size to 10 for all the datasets. STAN designs a spatiotemporal attention network capturing interactions between non-adjacent locations and non-consecutive check-ins using spatiotemporal contexts. We set the embedding size, the maximum length for trajectory sequence, and the number of negative samples to {50, 100, 10}, respectively. GetNext exploits the user flow trajectory graph and time-aware category context embedding for next location prediction; for a fair comparison, we do not use category information by assigning the same category to all POIs. We set the embedding size of POIs and users, the numbers of three GCN layers’ channels, the embedding size of the transformer encoder to {128, 32, 64, 128, 1024}, respectively. Flashback aggregates past hidden states with high predictive power using a context-aware weighting mechanism; Graph-Flashback incorporates a User-POI knowledge graph into the Flashback framework, where the social relationships between users are also used. We set $\alpha$ and $\beta$ as suggested by Flashback in Graph-Flashback and Flashback.
Appendix C Experiment Settings
We developed our model using the PyTorch framework and conducted experiments on the following hardware platform (CPU: Intel(R) Xeon(R) Gold 5320, GPU: NVIDIA GeForce RTX 3090). We empirically set the dimension of hidden states and all (POI, timestamp, and user) embedding sizes as 10, and the temporal decay factor $\alpha$ and spatial decay factor $\beta$ following the suggested settings in Flashback.
Table 6: Coefficient of variation in MRR over different RNN architectures (RNN, LSTM, GRU). A large value indicates a large variation in the performance over different RNNs.
@@ Appendix C Experiment Settings
Appendix D Ablation Study
We conduct an ablation study on our proposed method REPLAY, considering the following variants.
For each method, we instantiate it with the three common RNN architectures (vanilla RNN, LSTM, or GRU). Table 5 shows the results. First, we observe that our smoothed timestamp embeddings can significantly improve the location prediction performance by flexibly capturing the time-varying temporal regularities of human mobility. The improvement is evidenced by the superiority of REPLAY over REPLAY-noSTE (20.6% and 7.5% improvement on Gowalla and Foursquare, respectively). Second, we observe that the prediction conditioned on query time can also improve the location prediction performance as the query time intuitively provides additional information for location prediction problems, which results in the improvements of REPLAY over REPLAY-noQT (18.0% and 11.1% improvement on Gowalla and Foursquare, respectively). Finally, compared to Flashback, REPLAY integrating smoothed timestamp embeddings and making predictions conditioned on query time yields a significant improvement of 43.9% and 21.8% on Gowalla and Foursquare, respectively.
To future explore the impact of different granularities and the timestamp-specific learnable bandwidths, we conducted experiments on two variants of REPLAY, named REPLAY-MultiG (Multi-Granularity) and REPLAY-FixedB (Fixed Bandwidth). From Table 5, we observe that REPLAY consistently yields better performance (3.4% and 0.6% improvement on Gowalla and Foursquare, respectively) over REPLAY-MultiG, which further validates our design choice of hour-in-week timestamps. Besides, we observe that REPLAY with timestamp-specific learnable bandwidths achieves consistently better performance (1.5% and 0.6% improvement on Gowalla and Foursquare, respectively) over REPLAY-FixedB, which shows the necessity of using timestamp-specific learnable bandwidths in location prediction.
In addition, we also investigate the performance of using different RNN architectures (vanilla RNN, LSTM, or GRU). We observe that the vanilla RNN yields the best performance in general, while LSTM and GRU show comparable results. This is probably due to the fact that the flashback mechanism is able to effectively find useful historical hidden states using spatiotemporal contexts for location prediction, which indeed weakens the utility of retaining long-term memory by LSTM or GRU.
Interestingly, we find that REPLAY is more robust than Flashback against different RNNs. Specifically, we measure the robustness using a coefficient of variation, defined as the ratio of the standard deviation to the mean of a variable. We compute the coefficient of variation on the performance using different RNNs, and compare the results from different methods. Table 6 shows the results. We observe that compared to basic RNNs, Flashback can effectively reduce the coefficient of variation from 21.73% and 41.84% to 5.59% and 2.55% on Gowalla and Foursquare, respectively. More importantly, REPLAY can further reduce the coefficient of variation to 0.73% and 1.26% on Gowalla and Foursquare, respectively. In other words, integrating smoothed timestamp embeddings and making predictions conditioned on query time can make our model more robust against different RNNs.
(a) Hours in week
@@ Appendix D Ablation Study
Figure 5: The learnt bandwidths and the corresponding location prediction performance across different timestamps on Gowalla.
Appendix E The Time-Varying Temporal Regularities on Gowalla
As shown in Figure 5, similar to the Foursquare dataset, we observe a clear daily pattern, where the bandwidths in the morning have smaller values in general compared to other time periods and the bandwidths on weekends have larger values than on weekdays. Despite the daily pattern of the learnt bandwidths, we also observe a clear daily pattern (an inverse trend compared to the bandwidths) in location prediction performance.

### v3

@@ 4.2 Location Prediction Performance
Table II shows overall location prediction performance. First, we find that REPLAY achieves the best performance in most cases, yielding 7.7% and 10.9% improvement over the best-performing baselines on Gowalla and Foursquare, respectively. Moreover, REPLAY significantly outperforms the best baselines in Acc@1 (by 23.4% and 25.8% on Gowalla and Foursquare, respectively). The only exception is on Acc@10 where Graph-Flashback achieves slightly better performance than REPLAY; note that beyond user mobility trajectories, Graph-Flashback further uses user social relationships as input.
@@
We also highlight in Table II the methods that make predictions using query time information in different formats (e.g., a query timestamp, a query time interval, or a query time threshold). However, we did not observe a consistent improvement of these methods over others that make predictions without using query time information. This is due to the fact that the distinct models adopted by these baseline methods dominate the prediction performance. Moreover, the query time information is often used by existing methods on an ad-hoc basis without a systematic study of its impact. In this context, we provide a systematic ablation study below to verify the usefulness of the query time in location prediction.
TABLE III: Ablation study on REPLAY
Method Gowalla Foursquare
Acc@1 Acc@5 Acc@10 MRR Acc@1 Acc@5 Acc@10 MRR
Flashback (RNN) 0.1158 0.2754 0.3479 0.1925 0.2496 0.5399 0.6236 0.3805
Flashback (LSTM) 0.1024 0.2575 0.3317 0.1778 0.2398 0.5169 0.6014 0.3654
Flashback (GRU) 0.0979 0.2526 0.3267 0.1731 0.2375 0.5154 0.6003 0.3631
REPLAY-noSTE (RNN) 0.1362 0.3099 0.3845 0.2185 0.3055 0.5690 0.6415 0.4251
REPLAY-noSTE (LSTM) 0.1263 0.2980 0.3724 0.2079 0.3002 0.5663 0.6396 0.4209
REPLAY-noSTE (GRU) 0.1356 0.3099 0.3856 0.2183 0.2993 0.5634 0.6360 0.4192
REPLAY-noQT (RNN) 0.1397 0.3169 0.3924 0.2237 0.2817 0.5744 0.6489 0.4139
REPLAY-noQT (LSTM) 0.1343 0.3123 0.3873 0.2187 0.2739 0.5620 0.6360 0.4043
REPLAY-noQT (GRU) 0.1319 0.3105 0.3866 0.2168 0.2716 0.5623 0.6376 0.4029
REPLAY-MultiG (RNN) 0.1856 0.3357 0.3942 0.2570 0.3501 0.5894 0.6578 0.4598
REPLAY-MultiG (LSTM) 0.1808 0.3314 0.3904 0.2527 0.3440 0.5816 0.6493 0.4529
REPLAY-MultiG (GRU) 0.1798 0.3310 0.3897 0.2520 0.3426 0.5797 0.6477 0.4513
REPLAY-FixedB(RNN) 0.1861 0.3419 0.4024 0.2601 0.3513 0.5899 0.6575 0.4607
REPLAY-FixedB (LSTM) 0.1837 0.3411 0.4032 0.2587 0.3438 0.5817 0.6498 0.4528
REPLAY-FixedB (GRU) 0.1817 0.3374 0.3991 0.2558 0.3418 0.5807 0.6487 0.4512
REPLAY (RNN) 0.1866 0.3476 0.4124 0.2635 0.3529 0.5953 0.6648 0.4638
REPLAY (LSTM) 0.1848 0.3467 0.4100 0.2618 0.3448 0.5850 0.6544 0.4549
REPLAY (GRU) 0.1826 0.3449 0.4091 0.2597 0.3430 0.5831 0.6524 0.4530
4.3 Ablation Study
We consider the following variation of REPLAY and results are shown in Table III.
For each method, we instantiate it with the three common RNN architectures (vanilla RNN, LSTM, or GRU). First, we notice that REPLAY outperforms REPLAY-noSTE by 20.6% and 7.5% on Gowalla and Foursquare, respectively. This demonstrates our smoothed timestamp embeddings indeed boost performance significantly by flexibly capturing the time-varying temporal regularities of human mobility. Second, we find that REPLAY outperforms REPLAY-noQT by 18.0% and 11.1% on Gowalla and Foursquare, respectively, which shows the effectiveness of our prediction conditioned on query time as the query time intuitively provides additional information for location prediction problems. Moreover, REPLAY significantly outperforms our previous work Flashback by 43.9% and 21.8% on Gowalla and Foursquare, respectively, which verifies the effectiveness of integration smoothed timestamp embeddings and making predictions conditioned on query time.
To future explore the impact of different granularities and the timestamp-specific learnable bandwidths, we conducted experiments on two variants of REPLAY, named REPLAY-MultiG (Multi-Granularity) and REPLAY-FixedB (Fixed Bandwidth). From Table III, we observe that REPLAY consistently yields better performance (3.4% and 0.6% improvement on Gowalla and Foursquare, respectively) over REPLAY-MultiG, which further validates our design choice of hour-in-week timestamps. Besides, we find that REPLAY with timestamp-specific learnable bandwidths consistently outperforms, with improvements of 1.5% on Gowalla and 0.6% on Foursquare compared to REPLAY-FixedB. This highlights the importance of utilizing timestamp-specific learnable bandwidths in location prediction.
In addition, we also investigate the performance of using different RNN architectures (vanilla RNN, GRU or LSTM). We observe that the vanilla RNN yields the best performance in general, while LSTM and GRU show comparable results. This is probably due to the fact that the flashback mechanism is able to effectively find useful historical hidden states using spatiotemporal contexts for location prediction, which indeed weakens the utility of retaining long-term memory by LSTM or GRU.
Interestingly, we find that REPLAY is more robust than Flashback against different RNNs. Specifically, we use the coefficient of variation [67] to evaluate robustness, which is the standard deviation ratio to a variable’s mean. We compute the coefficient of variation on the performance using different RNNs, and compare the results from different methods. Table IV shows that compared to basic RNNs, Flashback can effectively reduce the coefficient of variation from 21.73% and 41.84% to 5.59% and 2.55% on Gowalla and Foursquare, respectively. More importantly, REPLAY can further reduce the coefficient of variation to 0.73% and 1.26% on Gowalla and Foursquare, respectively. In other words, integrating smoothed timestamp embeddings and making predictions conditioned on query time can make our model more robust against different RNNs.
TABLE IV: Coefficient of variation in MRR over Different RNN Architectures (RNN, GRU, LSTM).
@@ 4.3 Ablation Study
TABLE V: Settings on Time Granularities and Scales
Scale Granularity #Timestamps
Day Minute 1,440
Hour 24
Weekday& Weekend Minute 2,880
Hour 48
Week Minute 10,080
Hour 168
TABLE VI: Impact of different time granularities and scales on location prediction performance
Scale Granularity Gowalla Foursquare
Acc@1 Acc@5 Acc@10 MRR
Time per
epoch
Acc@1 Acc@5 Acc@10 MRR
Time per
epoch
Day Minute 0.1278 0.2918 0.3610 0.2060 41.35s 0.2801 0.5723 0.6468 0.4121 122.14s
Hour 0.1777 0.3411 0.4063 0.2555 39.35s 0.3483 0.5927 0.6622 0.4601 101.85s
Weekday& Weekend Minute 0.1804 0.3423 0.4090 0.2577 44.98s 0.3530 0.5938 0.6631 0.4633 142.74s
Hour 0.1821 0.3442 0.4097 0.2594 37.69s 0.3505 0.5941 0.6632 0.4618 102.33s
Week Minute 0.1756 0.3276 0.3891 0.2485 72.41s 0.3532 0.5940 0.6638 0.4636 264.00s
Hour 0.1866 0.3476 0.4124 0.2635 44.20s 0.3529 0.5953 0.6648 0.4638 103.49s
4.4 Temporal Regularities on Different Time Granularities and Scales
We investigate the impact of different time granularities (i.e., minute or hour) and time scales (i.e., day, weekday&weekend or week) in our smoothed timestamp embeddings used by REPLAY. Table V shows the settings we considered and their corresponding number of timestamps. Note that for the time scale weekday&weekend, we define two sets of daily cyclical timestamps for weekday and weekend, respectively; the similarity between the timestamps in each set is computed independently from the other set. This setting is motivated by the fact that human mobility patterns are often split into weekday and weekend patterns by existing studies [34]. The performance comparison between different settings is shown in Table VI.
First, we observe that the week scale is consistently better than the day scale. For example, on the hour granularity (the best time granularity setting as we discuss below), the week scale outperforms the day scale by 2.89% and 0.74% on Gowalla and Foursquare, respectively. This is due to the fact that the former can capture the different temporal regularities across different days in a week, in particular the different regularities on weekdays and weekends (more detail in Section 4.5 below), while the latter cannot. Moreover, compared to the weekday&weekend scale which can also capture the different regularities on weekdays and weekends, the week scale still achieves consistently better performance (with 1.43% and 0.4% improvement on Gowalla and Foursquare, respectively, under the hour granularity). Because compared to the weekday&weekend scale, the week scale can subtly capture the regularity differences between different weekdays (also between different days in the weekend).
Second, we observe that the hour granularity is generally better than the minute granularity, as the minute granularity is often over-specific to model the temporal regularity of user mobility. One exception is on the week scale on Foursquare where we observe comparable results; however, in this case, the minute granularity shows significantly worse runtime performance (more than twice the training time as shown in Table VI) than the hour granularity, as it needs to learn much more embeddings for minute-in-week timestamps (10,080 in total) than for hour-in-week timestamps (168 in total). Moreover, we also conducted experiments on multi-granularity combining both hour-in-day and day-in-week timestamp embeddings in Section 4.3; the results show 3.4% and 0.6% improvement of our REPLAY over this multi-granularity setting, on Gowalla and Foursquare, respectively. In summary, we choose hour-in-week timestamps in REPLAY as the default setting.
(a) Hours in a week
@@ 4.4 Temporal Regularities on Different Time Granularities and Scales
Figure 4: The learnt bandwidths and the corresponding location prediction performance across different timestamps.
4.5 Revealing the Time-Varying Temporal Regularities
In this experiment, we investigate the time-vary temporal regularities of user mobility using the learnt bandwidths from REPLAY. Figure 4(a) shows the values of the learnt bandwidths for the 168 hour-in-week timestamps and the corresponding location prediction performance in MRR for each timestamp on both Gowalla and Foursquare datasets.
First, we find a clear daily pattern, where the bandwidths in the morning have smaller values in general compared to other time periods; this implies a stronger temporal regularity of the morning mobility (often working-related activities), as the smoothed timestamp embeddings require less information from the neighboring timestamps for location prediction. In contrast, the learnt bandwidths in the nighttime often have larger values, implying a weaker temporal regularity in the nighttime (often entertainment-related activities in LBSNs), as the smoothed timestamp embeddings require more information from the neighboring timestamps for location prediction. Moreover, we further compare the bandwidths between weekdays and weekends, by plotting the average bandwidth for each hour on weekdays and on weekends, as shown in Figure 4(b) and 4(c), respectively. We observe that despite the similarity of having smaller bandwidths in the morning, the bandwidths on weekends have larger values than on weekdays, which implies that weekend activities are less regular than weekdays in general. Moreover, we find that the weekend afternoon sometimes shows the weakest regularity evidenced by the largest learnt bandwidth.
Second, we also discover a distinct daily pattern in the performance of location prediction. As shown in Figure 4, the location prediction performance shows an inverse trend compared to the bandwidths. This is due to the fact that the weaker regularity of user mobility at a certain timestamp (in the nighttime or on weekends, for example) makes it harder for location prediction; subsequently, the learnt bandwidth has a larger value, as the query timestamp is less confident for location prediction and thus resorts more to its neighboring timestamps for prediction, and vice versa.
TABLE VII: Location prediction performance and the corresponding average bandwidth in different prediction periods.
Prediction time period Gowalla Foursquare
Average MRR Average Bandwidth Average MRR Average Bandwidth
Daytime (6:00-18:00) 0.2767 0.85 0.4856 1.76
Nighttime (18:00-6:00) 0.2362 0.88 0.4302 2.00
Weekday 0.2922 0.82 0.4841 1.84
Weekend 0.1898 0.98 0.3758 2.00
Finally, we quantitatively verify our motivating example where the strength of the temporal regularities varies across typical time periods (daytime/nighttime and weekday/weekend). Table VII shows the average MRR and learnt bandwidths in typical time periods. We see that the daytime prediction performance is 15.0% better (with 7.5% lower bandwidth) on average than the nighttime, while the weekday prediction performance is 41.4% better (with 11.9% lower bandwidth) on average than the weekend.
5 Conclusion
In this paper, motivated by the time-varying temporal regularities of user mobility, we propose REPLAY, a general RNN architecture designed to capture such varying regularities for location prediction. REPLAY seamlessly incorporates smoothed timestamp embeddings with learnable bandwidths into a flashback mechanism. The timestamp-specific learnable bandwidths can automatically learn to adapt to the temporal regularities of different strengths across different timestamps. We conduct a thorough evaluation using two real-world LBSN datasets. Results show that REPLAY outperforms a wide range of state-of-the-art location prediction methods, with 7.7%-10.9% improvement over the best-performing baselines. Moreover, we reveal interesting temporal regularity patterns using the learnt bandwidths: 1) morning mobility consistently shows a stronger regularity compared to other time periods; 2) weekend mobility is less regular than weekdays in general, while the weekend afternoon sometimes shows the weakest regularity.
In the future, we plan to further investigate the temporal regularities varying across different user groups and POI categories.